OpenStudy (anonymous):

The function f(t) = 65 sin (pi over 5t) + 35 models the temperature of a periodic chemical reaction where t represents time in hours. What are the maximum and minimum temperatures of the reaction, and how long does the entire cycle take? Maximum: 65°; minimum: 35°; period: pi over 5 hours Maximum: 100°; minimum: −30°; period: pi over 5 hours Maximum: 65°; minimum: 35°; period: 10 hours Maximum: 100°; minimum: −30°; period: 10 hours

OpenStudy (anonymous):

which one would you guess?

OpenStudy (amilapsn):

did you mean \[f(t)=65\sin(\frac{ \pi }{ 5 }t) +35\]?

OpenStudy (amilapsn):

If you've asked a trigonometric function like this all you have to know is \[-1 \le sinx \le +1\] for any rational x.... i.e. \[x\]

OpenStudy (amilapsn):

correction:\[x \epsilon \mathbb{R}\]